OpenStudy (anonymous):

Does i^2=sq root of -1?

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

does ok mean yes

OpenStudy (anonymous):

OK means "Do you understand, get it, agree etc" ie i^2 = -1 (i^2 does NOT = sq rt of -1)

OpenStudy (anonymous):

no- I don't get it

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

isn't sq. rt of anything always a positive number

OpenStudy (anonymous):

how do you get the square root of a negative number?

OpenStudy (anonymous):

sq rt of -25 is 5 or not

OpenStudy (anonymous):

How can it be? 5*5 = 25 and -5*-5 is also 25

OpenStudy (anonymous):

so what is sq rt of -25

OpenStudy (anonymous):

You know the rule sqrt (ab) = sqrt (a) sqrt(b) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Right, so we have sqrt(-25) = sqrt(25)sqrt(-1) Agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now we have i^2 = -1 and so we can think of i = sqrt(-1)

OpenStudy (anonymous):

ok I^2 =1 not -1 thank you

OpenStudy (anonymous):

So sqrt of -25 = 5i

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

5i*5i = 25i^2 = 25*-1 = -25

OpenStudy (anonymous):

so now I am lost again I thought i^2=1 but no it =-1

OpenStudy (anonymous):

I keep telling u i^2 = -1 (it's a definition not something you have to prove)

OpenStudy (anonymous):

ok I was missing that. I thought it wa equal to 1

OpenStudy (anonymous):

U see how it works now?

OpenStudy (anonymous):

Strictly speaking, if i^2 = -1 then there should be 2 roots i and -i but by convention we pick i as the principal one (similar to positive sq rt convention)

OpenStudy (anonymous):

yes, and I am looking in my book at the -sq.rt. and what you are saying makes sense now. Thanks for the lesson

OpenStudy (anonymous):

ur welcome